Complete the following data taken from the condensed income statements for merchandising Companies X, Y, and Z. Enter net loss with a minus sign.

Company X Company Y Company Z
Net income/(net loss) $220 $ $(40)
Sales 1,260 895
Gross profit 400 480
Operating expenses 540
Cost of goods sold 310 750
Answer #1

For Company X

Gross profit = Sales - Cost of goods sold

400 = Sales - 310

Sales = 710

Net income = Gross profit- Operating expenses

220 = 400-Operating expenses

Operating expenses = 400-220

= $180

For Company Y

Gross profit = Sales - Cost of goods sold

= 1,260-750

= $510

Net income/Loss = Gross profit- Operating expenses

= 510-540

= ($30)

Company C

Gross profit = Sales - Cost of goods sold

480 = 895-Cost of goods sold

Cost of goods sold = 895-480

= $415

Net income/Loss = Gross profit- Operating expenses

-40 = 480- Operating expenses

Operating expenses = 480+40

= $520

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks